OpenStudy (anonymous):

Sawyer spent $45.22 on short-sleeve and long-sleeve T-shirts. The long-sleeve T-shirts cost $13.99 each. If he bought two of each type of shirt, what was the price of each short-sleeve T-shirt?

OpenStudy (anonymous):

45.22=13.99(2)+2x

OpenStudy (anonymous):

Solve this and you'll get the answer

OpenStudy (anonymous):

45.22=27.98+2x -27.98 -27.98 17.24=2x divide each by 2 8.62=x